OpenStudy (anonymous):

Harvey is 3 times as old as Jane. The sum of their ages is 48 years. Find the ages of Harvey and Jane.

OpenStudy (johnweldon1993):

H = 3j H + J = 48 we know what H equals so sub it into the second equation 3J + J = 48 4J = 48 J = 12 so If Jane is 12 Check to see if H = 3j H = 3(12) H = 36 does 36 + 12 = 48? yes and is 36, 3 times more than 12? yes it is

OpenStudy (anonymous):

yes it is